# A Note on Sample Complexity of Interactive Imitation Learning with Log Loss

### v1

@@ B.2 New Guarantees for DAgger with Trajectory-wise Annotation
Then, Lemma 22 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s_{1}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(s^{n}_{1:H}),\pi^{\mathrm{E}}(s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
Finally, the second part of the theorem follows by applying Lemma 14.
Appendix C DAgger Variant with State-wise Annotation
Different from [4], instead of having access to action demonstrations at each step $h$, at round $n$, we only sample one state $(s^{n},h^{n})$ uniformly from a trajectory ($d^{\pi^{n}}$) and query for expert’s annotation.
We recall the state-wise Hellinger estimation error
$\operatorname{OnEst}^{\mathrm{State}}_{N}=\sum_{n=1}^{N}\mathbb{E}_{(s,h)\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right].$
In the following, we show that performance difference between $\hat{\pi}$ returned from Algorithm 1 and $\pi^{\mathrm{E}}$ can also be bounded by $\operatorname{OnEst}^{\mathrm{State}}_{N}$ in Lemma 16. Then, we show the state-wise annotation complexity of Algorithm 1 in Theorem 17.
Lemma 16.
For any MDP $\mathcal{M}$, deterministic expert $\pi^{\mathrm{E}}$, and sequence of policies $\left\{\pi^{n}\right\}_{n=1}^{N}$, each of which can be Markovian or a first-step mixture of Markovian policies, their first step uniform mixture policy $\hat{\pi}$ satisfies.
Proof.
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu\cdot\lambda(\hat{\pi}\parallel\pi^{\mathrm{E}}),$
where we recall the notation that
$\lambda(\pi\parallel\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}\mathbb{I}(a_{h}\neq\pi^{\mathrm{E}}_{h}(s_{h}))\right]=\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi_{h}(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
The proof follows by upper-bounding $\sum_{n=1}^{N}\lambda(\pi^{n}\parallel\pi^{\mathrm{E}})$ by $H\cdot\operatorname{OnEst}^{\mathrm{State}}_{N}$. To this end, it suffices to show that for any Markovian policy $\pi$,
$H\cdot\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D_{H}^{2}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right]\geq\frac{1}{2}\sum_{h=1}^{H}\mathbb{E}^{\pi}\|\pi_{h}(s_{h})-\pi^{\mathrm{E}}(s_{h})\|_{1}.$
Observe that $H\cdot\mathbb{E}_{(s,h)\sim d^{\pi}}\left[D_{H}^{2}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\right]=\sum_{h=1}^{H}\mathbb{E}^{\pi}\left[D_{H}^{2}(\pi_{h}(s_{h}),\pi^{\mathrm{E}}_{h}(s_{h}))\right]$, we conclude the proof by applying Lemma 18 with $p=\pi_{h}(s_{h})$ and $q=\pi^{\mathrm{E}}_{h}(s_{h})$, which gives
$D^{2}_{\text{H}}(\pi_{h}(s),\pi^{\mathrm{E}}_{h}(s))\geq\frac{1}{2}\|\pi_{h}(s)-\pi^{\mathrm{E}}_{h}(s)\|_{1}.$
Theorem 17 (Theorem 3 Restated).
If Algorithm 1 is run with a deterministic expert policy $\pi^{E}$, an MDP $\mathcal{M}$ such that $(\mathcal{M},\pi^{E})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that realizability holds, and the online learning oracle $\mathbb{A}$ set to exponential weight algorithm (see Proposition 21). Then, with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N}\leq\log(B)+2\log(1/\delta),$
and furthermore, the returned $\hat{\pi}$ satisfies
$J(\hat{\pi})-J(\pi^{\mathrm{E}})\leq\mu H\frac{\log(B)+2\log(1/\delta)}{N}.$
Proof.
Recall the each-step mixing in Definition 2, since $\bar{\pi}_{u}$ is a each-step mixing policy, where $\forall h\in[H],s\in{\mathcal{S}}$, $\bar{\pi}_{u,h}(a|s)=\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a|s)$.
By using $\bar{\pi}_{u}$, the loss functions at each round that passed through $\textbf{Alg}_{\text{Est}}$, $\ell^{n}(u)$ is of the form
$\ell^{n}(u)=\log(1/\bar{\pi}_{u,h}(a^{n,*}|s^{n}))=\log\left(\frac{1}{\sum_{\pi\in\mathcal{B}}u(\pi)\pi_{h}(a^{n,*}|s^{n})}\right),$
which is 1-exp-concave with respect to $u$. Thus, implementing $\textbf{Alg}_{\text{Est}}$ using the exponential weights algorithm (Proposition 21) achieves:
$\sum_{n=1}^{N}\log(1/\pi^{n}_{h^{n}}(a^{*,n}\mid s^{n}))\leq\sum_{n=1}^{N}\log(1/\pi^{\mathrm{E}}_{h^{n}}(a^{*,n}\mid s^{n}))+\log(B)=\log(B).$
Then, Lemma 22 with $x^{n}=(s^{n},h^{n})$, $y^{n}=a^{*,n}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(s^{n},h^{n},a^{n},a^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{State}}_{N}=\sum_{n=1}^{N}\mathbb{E}_{(s^{n},h^{n})\sim d^{\pi^{n}}}\left[D^{2}_{\text{H}}(\pi^{n}_{h^{n}}(s^{n}),\pi^{\mathrm{E}}_{h^{n}}(s^{n}))\right]\leq\log(B)+2\log(1/\delta).$
The second part of the theorem follows by applying Lemma 16.
Appendix D Auxiliary Results
@@
Lemma 19 (Performance Difference Lemma [7, 13]).
For two Markovian policies $\pi$ and $\pi^{\mathrm{E}}$ $:{\mathcal{S}}\to\Delta(\mathcal{A})$, we have
$J(\pi)-J(\pi^{\mathrm{E}})=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}A_{h}^{E}(s_{h},a_{h})\right],$
@@
Lemma 20 (Lemma D.2. of  [4]).
@@
Proposition 21 (Proposition 3.1 of [2]).
@@
$\sum_{n=1}^{N}\ell^{n}(u^{n})\leq\min_{x\in\mathcal{X}}\sum_{n=1}^{N}\ell^{n}(e_{x})+\frac{\ln|\mathcal{X}|}{\eta}.$
Lemma 22 (Restatement of Lemma A.14 in [5]).
@@
We have the following well-known lemma for causally-conditioned probability (e.g. [21]).
@@
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})$
@@
and for any first-step mixture of Markovian policy $\pi_{u}$,
@@
$P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
@@
$\sum_{\nu}u(\nu)P_{0}(s_{1})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$

### v2

@@ E.4 New Guarantees for DAgger Variant with Trajectory-wise Annotation
Then, Lemma 39 with $x^{n}=s^{n}_{1:H}$, $y^{n}=a^{*,n}_{1:H}$, $g_{*}=\pi^{\mathrm{E}}$, and $\mathcal{H}^{n}=\{o^{n^{\prime}}\}_{n^{\prime}=1}^{n}$, where $o^{n}=\left(\mathbf{E^{\prime}}^{n},a_{1}^{n},a_{1}^{*,n},\dots,s_{H}^{n},a_{H}^{n},a_{H}^{*,n}\right)$, implies that with probability at least $1-\delta$,
$\operatorname{OnEst}^{\mathrm{Traj}}_{N}=\sum_{n=1}^{N}\mathbb{E}^{\pi^{n}}\left[D^{2}_{\text{H}}(\pi^{n}(\cdot\mid s^{n}_{1:H}),\pi^{\mathrm{E}}(\cdot\mid s^{n}_{1:H}))\right]\leq\log(B)+2\log(1/\delta).$
Finally, the second part of the theorem follows by applying Lemma 33. ∎
Appendix F Auxiliary Results
@@
Lemma 36 (Performance Difference Lemma [Kakade2002ApproximatelyOA][ross2010efficient]).
For two Markovian policies $\pi$ and $\pi^{\mathrm{E}}$ ${\mathcal{S}}\to\Delta(\mathcal{A})$, we have
$J(\pi^{\mathrm{E}})-J(\pi)=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}A_{h}^{E}(s_{h},a_{h})\right],$
@@
Lemma 37 (Lemma D.2. of  [foster2024behavior]).
@@
Proposition 38 (Proposition 3.1 of [cesa2006prediction]).
@@
$\sum_{n=1}^{N}\ell^{n}(u^{n})\leq\min_{x\in\mathcal{X}}\sum_{n=1}^{N}\ell^{n}(e_{x})+\frac{\log|\mathcal{X}|}{\eta}.$
Lemma 39 (Restatement of Lemma A.14 in [foster2021statistical]).
@@
We have the following well-known lemma for causally-conditioned probability (e.g. [ziebart2010modeling]).
@@
$\mathbb{P}^{\mathcal{M}}(s_{1:H}\parallel a_{1:H-1})=P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})$
@@
and for any first-step mixing of Markovian policy $\pi_{u}$,
@@
$P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\pi_{h}(a_{h}\mid s_{h}).$
@@
$\sum_{\nu}u(\nu)P_{0}(\mathbf{E^{\prime}})\prod_{h=1}^{H-1}P_{h}(s_{h+1}\mid s_{h},a_{h})\prod_{h=1}^{H}\nu_{h}(a_{h}\mid s_{h}).$
Appendix G Experiment Details
We compare Warm-Stagger with Behavior Cloning (BC) and Stagger on continuous-control tasks from OpenAI Gym MuJoCo [todorov2012mujoco, brockman2016openai] with episode length $H=1000$.
Infrastructure and Implementation.
All experiments were conducted on a Linux workstation equipped with an Intel Core i9 CPU (3.3GHz) and four NVIDIA GeForce RTX 2080 Ti GPUs. Our implementation builds on the publicly available DRIL framework [brantley2019disagreement] (https://github.com/xkianteb/dril), with modifications to support online learning. The continuous control environments used in our experiments are: “HalfCheetahBulletEnv-v0”, “AntBulletEnv-v0”, “Walker2DBulletEnv-v0”, and “HopperBulletEnv-v0”. We include an anonymous link to our implementation here: https://github.com/liyichen1998/Interactive-and-Hybrid-Imitation-Learning-Provably-Beating-Behavior-Cloning.
Environments and Expert Policies.
We use four MuJoCo environments: Ant, Hopper, HalfCheetah, and Walker2D. The expert policy is a deterministic MLP pretrained via TRPO [schulman2015trust, schulman2017proximal], with two hidden layers of size 64.
Learner Architecture.
The learner uses the same MLP architecture as the expert. Following [foster2024behavior], we use a diagonal Gaussian policy:
$\pi(a\mid s)=\mathcal{N}\left(f_{\theta}(s),\operatorname{diag}(\sigma^{2})\right),$
where $f_{\theta}(s)\in\mathbb{R}^{d_{\mathcal{A}}}$ is the learned mean, and $\sigma\in\mathbb{R}^{d_{\mathcal{A}}}$ is a learnable log-standard deviation vector.
Each model is trained from random initialization using a batch size of 100, a learning rate of $10^{-3}$, and up to 2000 passes over the dataset, with early stopping evaluated every 250 passes using a 20% held-out validation set.
Learning Protocols.
To evaluate the performance of BC against the number of states annotated, we reveal expert state-action pairs sequentially along expert trajectories until the annotation budget is reached. For Stagger, each round it rolls out the latest policy, samples a state uniformly from the trajectory, queries it for the expert action, and updates immediately.
For Warm-Stagger, we begin with BC and switch to Stagger after a predefined number of offline state-action pairs has been used, denoted as $N$. We set $N$ to be 100, 200, or 400 for easier tasks (e.g., Hopper, Ant) and 200, 400, or 800 for harder tasks (e.g., HalfCheetah, Walker2D).
Cost Model and Evaluation.
We assign a cost of 1 to each offline state-action pair and a cost of $C=1$ or $2$ to each interactive query. We run each method for 10 random seeds. For every 50 new state-action pairs collected, we evaluate the current policy by running 25 full-episode rollouts and reporting the average return.
Though the nonrealizable setting is beyond the scope of this work, we expect that some variant of our algorithm can still give reasonable performance, provided that the policy class is expressive enough (so that the approximation error is nonzero but small). For example, [li2023agnostic] observed that with nonrealizable stochastic experts, DAgger variants outperform BC, and exhibit learning curves similar to ours.
G.1 Additional Experiment Plots
We present extended experiment results with larger cost budgets. As shown in Figure 4, we allocate a total annotation budget of 2000 for Hopper and Ant, and 4000 for HalfCheetah and Walker. This complements Figure 3 in the main paper by showing the full training curves without zooming into the stage with small cost budget. The trends are consistent with our earlier observations: Warm-Stagger achieves similar or better sample efficiency compared to Stagger when $C=1$, and clearly outperforms both baselines under the cost-aware setting where $C=2$.
Figure 4: Sample and cost efficiency on MuJoCo tasks. The top row shows expected return vs. number of annotations ($C=1$); the bottom row shows performance under a cost-aware setting ($C=2$). Warm-Stagger (WS) is initialized with 1/20, 1/10, or 1/5 of the samples as offline demonstrations. It matches Stagger in sample efficiency and outperforms the baselines when $C=2$, especially WS(1/5).
G.2 Experiment with MSE Loss
We additionally evaluate our algorithms using mean squared error (MSE) as the loss function for optimization. All training settings remain identical to the main experiments with log loss, except that we use a learning rate of $2.5\times 10^{-4}$. As shown in Figure 5, we observe qualitatively similar results to those under log loss shown in 3, consistent with prior observations in [foster2024behavior], with the added benefit of more stable training dynamics.
Figure 5: Performance comparison under MSE loss across MuJoCo tasks. Results show that Warm-Stagger (WS) achieves comparable sample efficiency and performance to the log loss setting, with improved training stability. Each curve represents the average over 10 seeds.
G.3 Additional Experiments with Algorithm 3
For completeness, we evaluate Tragger and its warm-start variant (Warm-Tragger), as shown in Algorithm 4, on continuous control tasks and the same MDP setup as in Figure 2. The key distinction between Warm-Tragger and Warm-Stagger lies in the annotation mode: the former employs trajectory-wise oracle feedback instead of state-wise annotation, leading to notably different behaviors, as shown in Figure 6.
In particular, for Ant and HalfCheetah, the sample efficiency ($C=1$) of Tragger and Warm-Tragger is significantly worse than that of Stagger due to the cold-start problem: early DAgger rollouts have poor state coverage but must still proceed until the end of each trajectory. In contrast, Stagger samples only a single state per trajectory, thereby better leveraging interactive feedback.
For Hopper and Walker, however, Tragger and Warm-Tragger achieve performance closer to Stagger. These environments feature hard resets when the agent fails (unlikely in Ant and never in HalfCheetah), which truncate poor trajectories and consequently improve sample efficiency.
Overall, these observations suggest a natural middle ground between full-trajectory and single-state annotation—namely, batch queries (e.g., sampling 50 states per trajectory), as explored by [li2023agnostic] with comparable results.
A head-to-head comparison between Tragger and Stagger, as well as between Warm-Tragger and Warm-Stagger, is shown in Figure 7, highlighting the advantage of state-wise over trajectory-wise annotation.
However, this advantage does not hold in general: in the toy MDP in Figure 2, Tragger and Warm-Tragger achieve performance nearly identical to Stagger and Warm-Stagger, as shown in Figure 8.
Algorithm 4 Warm-Tragger: Warm-start Tragger with offline demonstrations
1:Input: MDP $\mathcal{M}$, trajectory-wise expert annotation oracle $\mathcal{O}^{\mathrm{Traj}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$, offline expert dataset $D_{\text{off}}$ of size $N_{\mathrm{off}}$, online budget $N_{\mathrm{int}}$
2: Initialize $\mathbb{A}$ with policy class $\mathcal{B}_{\text{bc}}:=\{\pi\in\mathcal{B}:\pi(s)=\pi^{\mathrm{E}}(s)\ ,\forall s\in D_{\text{off}}\}$.
3:for $n=1,\dots,N_{\mathrm{int}}/H$ do
4:  Query $\mathbb{A}$ and receive $\pi^{n}$.
5:  Execute $\pi^{n}$ and sample $s^{n}_{1:H}$ following $\mathbb{P}^{\pi^{n}}$. Query $\mathcal{O}^{\mathrm{Traj}}$ for $a_{1:H}^{*,n}=\pi^{\mathrm{E}}(s^{n}_{1:H})$.
6:  Update $\mathbb{A}$ with loss function
$\ell^{n}(\pi):=\log\left(\frac{1}{\pi(a_{1:H}^{*,n}\parallel s_{1:H}^{n})}\right).$ (23)
7:end for
8:Output: $\hat{\pi}$, a first-step uniform mixture of $\{\pi^{1},\dots,\pi^{N}\}$.
Figure 6: Sample efficiency of algorithms on MuJoCo tasks, showing expected return vs. number of annotations ($C=1$). Warm-Tragger (WT) is initialized with 1/20, 1/10, or 1/5 of the total annotation budget as offline demonstrations. Specifically, WT($n$) refers to WT with offline demonstrations of total length $n$. Although the performance of WT improves with more offline demonstrations, both Tragger and Warm-Tragger remain inferior to Stagger and, in many cases, even underperform Behavior Cloning, confirming the advantage of state-wise over trajectory-wise annotations.
Figure 7: Head-to-head sample efficiency comparison between Tragger and Stagger, and between Warm-Tragger and Warm-Stagger under equal (since we are talking about comparison here) offline demonstration budgets. Stagger and Warm-Stagger consistently outperform Tragger and Warm-Tragger. The performance gap narrows as the offline budget increases, effectively alleviating the cold-start problem suffered by Tragger.
Figure 8: Similar to Figure 2, we evaluate Tragger and Warm-Tragger (WT) with 200, 800, 3200 offline (state, expert action) pairs in the toy MDP therein. All methods are evaluated under equal total annotation cost with $C=1$. With 800 offline (state, expert action) pairs, WT significantly improves the sample efficiency over the baselines and explores $\mathbf{E^{\prime}}$ more effectively. The performance of Tragger and Warm-Tragger is almost the same as Stagger and Warm-Stagger in Figure 2.